# Nonequilibrium thermodynamics of populations of weakly-coupled low-temperature-differential Stirling engines with synchronous and asynchronous transitions

### v1

@@ III.5 Discussion
The dynamic behavior of the differential equation given in Eq. (58) has already been well investigated [13, 14, 15]. For $K<P_{0}/2$, no fixed point exists and the phase difference approaches a stable limit cycle circling the phase space $\mathbb{T}\times\mathbb{R}$. For $K>P_{0}/2$, there is a stable fixed point and a saddle point, which are created by a saddle-node bifurcation at $K=P_{0}/2$. The stable limit cycle coexists with the fixed points if the increase in kinetic energy due to $P_{0}$ compensates the decrease due to friction. Thus, the parameter space of the system can be divided into three regions: a region where a limit cycle is globally stable, a region where a fixed point is globally stable, and a region where both coexist. At the boundary that separates the globally stable fixed point regime and the coexistence regime, a homoclinic orbit of the saddle fixed point emerges from the limit cycle in a homoclinic bifurcation. This boundary can be obtained in the low-friction limit, which is given by $P_{0}\approx\frac{4\sqrt{2}}{\pi}\cdot\alpha\sqrt{K}$ and is found to be consistent with the actual value when $\alpha/\sqrt{K}<0.6$. The homoclinic boundary intersects the saddle-node bifurcation line ($K=P_{0}/2$) at a numerically determined value of $\alpha/\sqrt{K}\approx 1.69$. For $\alpha/\sqrt{K}\gtrsim 1.69$, the saddle-node bifurcation and the homoclinic bifurcation combine to a saddle-node homoclinic bifurcation.
@@
Comparing the original model (56a) and (56b) and the averaged differential equation (57a) and (57b), we find that the limit cycles corresponding to the stable and unstable $1:1$ synchronized states become stable and unstable fixed points, while limit cycles corresponding to $m:n$ synchronized states disappear through averaging, which implies that the devil’s staircase that appeared leading up to the synchronous transitions can not be observed for the conceptual model. This suggests that the model obtained by averaging fast variables preserves macroscopic structure but loses fine structure of the original model. The loss of the fine structure is the result of the fact that the dynamics of the differential system is completely separated from the 4-dimensional coupled system through the averaging method.
@@ V Summary and outlook
This study has presented a theory of quasilinear irreversible thermodynamics for multiple weakly-coupled LTD Stirling engines with synchronous and asynchronous transitions. We analyzed the effects of synchronous and asynchronous transitions on power and thermal efficiency using the quasilinear response relations between formally defined thermodynamic fluxes and forces, and constructed a conceptual model satisfying the quasilinear response relations to give a physical interpretation for the transitions in power and thermal efficiency. We also confirmed that the conceptual model can be obtained by averaging fast variables of the original model, and that unlike the quasilinear response relations, it preserved the irreversible nature in the relative motion of coupled Stirling engines. Furthermore, we compared the dynamics between the original and conceptual models for two-engine systems. Generalization of the theory developed in this study to a form applicable to a broader class of non-equilibrium autonomous heat engines is a challenging subject for future work.
@@ Acknowledgements
On the other hand, the averaged equation for the heat flux from the heat reservoir of temperature $T_{\mathrm{t}}$ is also given by Eq. (54), indicating that the sum of the heat fluxes from the two heat reservoirs is zero. Therefore, analysis using the energy conservation law requires the approximation of the heat flux to the second order for minute quantities. To this end, we reserve the temperature $T(\theta_{i},\omega_{i})$ to the second order for minute quantities instead of truncating $\mathcal{O}\left(\Delta T\omega_{i},\omega_{i}^{2}\right)$ as in Eq. (62):
@@
The first term in Eq. (82) coincides with the averaged brake power $\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{M}}\Bigr{\rangle}$ of the engine ${\mathrm{M}}_{\mathrm{EG}}$ in system M, which can be confirmed by multiplying both sides of Eq. (32) by $N\omega_{\mathrm{m}}$ and averaging it over time. On the other hand, the second term in Eq. (82) agrees with the averaged brake power $\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{D}}\Bigr{\rangle}$ for system D. Therefore, Eq. (82) agrees with the averaged brake power corresponding to the conceptual model constructed in III. B. It can be seen that part of the power dissipated by the viscous friction, $\sum_{i=1}^{N}\Gamma\Bigl{\langle}\omega_{i}^{2}\Bigr{\rangle}$, comes from the viscous friction due to the rotational motion of the Stirling engines, while the remaining, $\frac{\sigma^{2}}{G}\biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\biggr{\rangle}_{\theta}\left(\sum_{i=1}^{N}\Bigl{\langle}\omega_{i}^{2}\Bigr{\rangle}\right)$, comes from the loss of the energy received from the heat reservoirs. The latter is the heat leakage due to nonquasi-static processes associated with the rotational motion of the Stirling engines, which depends on the magnitude of the coupling strength: increasing the coupling strength results in smaller heat leakage. Thus, the conceptual model constructed in III. B interprets the heat leakage as energy dissipation due to viscous friction of the flywheels in system D, and the heat flux from the heat reservoir in system M is thus independent of the coupling strength.

### v2

@@ III.5 Discussion
The dynamic behavior of the differential equation given by Eq. (57) has already been well investigated [13, 14, 15]. For $K<P_{0}/2$, no fixed point exists and the phase difference approaches a stable limit cycle circling the phase space $\mathbb{T}\times\mathbb{R}$. For $K>P_{0}/2$, there is a stable fixed point and a saddle point, which are created by a saddle-node bifurcation at $K=P_{0}/2$. The stable limit cycle coexists with the fixed points if the increase in kinetic energy due to $P_{0}$ compensates the decrease due to friction. Thus, the parameter space of the system can be divided into three regions: a region where a limit cycle is globally stable, a region where a fixed point is globally stable, and a region where both coexist. At the boundary that separates the globally stable fixed point regime and the coexistence regime, a homoclinic orbit of the saddle fixed point emerges from the limit cycle in a homoclinic bifurcation. This boundary can be obtained in the low-friction limit, which is given by $P_{0}\approx\frac{4\sqrt{2}}{\pi}\cdot\alpha\sqrt{K}$ and is found to be consistent with the actual value when $\alpha/\sqrt{K}<0.6$. The homoclinic boundary intersects the saddle-node bifurcation line ($K=P_{0}/2$) at a numerically determined value of $\alpha/\sqrt{K}\approx 1.69$. For $\alpha/\sqrt{K}\gtrsim 1.69$, the saddle-node bifurcation and the homoclinic bifurcation combine to a saddle-node homoclinic bifurcation.
@@
Comparing the original model (55a) and (55b) and the averaged differential equation (56a) and (56b), we find that the limit cycles corresponding to the stable and unstable $1:1$ synchronized states become stable and unstable fixed points, while limit cycles corresponding to $m:n$ synchronized states disappear through averaging, which implies that the devil’s staircase that appeared leading up to the synchronous transitions can not be observed for the conceptual model. This suggests that the model obtained by averaging fast variables preserves the macroscopic structure but loses the detailed structure of the original model. The loss of the detailed structure is the result of the fact that the dynamics of the differential system is completely separated from the 4-dimensional coupled system through the averaging method.
@@ V Summary and outlook
This study has presented a theory of quasilinear irreversible thermodynamics for multiple weakly-coupled LTD Stirling engines with synchronous and asynchronous transitions. We analyzed the effects of synchronous and asynchronous transitions on power and thermal efficiency using the quasilinear response relations between formally defined thermodynamic fluxes and forces, and constructed a conceptual model satisfying the quasilinear response relations to give a physical interpretation for the transitions in power and thermal efficiency. We also confirmed that the conceptual model can be obtained by averaging fast variables of the original model, and that unlike the quasilinear response relations, it preserves the thermodynamic irreversibility of the coupled Stirling engines. Furthermore, we compared the dynamics between the original and conceptual models for two-engine systems. Generalization of the theory developed in this study to a form applicable to a broader class of non-equilibrium autonomous heat engines is a challenging subject for future work.
@@ Acknowledgements
On the other hand, the averaged equation for the heat flux from the heat reservoir of temperature $T_{\mathrm{t}}$ is also given by Eq. (53), indicating that the sum of the heat fluxes from the two heat reservoirs is zero. Therefore, analysis using the energy conservation law requires the approximation of the heat flux to the second order for minute quantities. To this end, we preserve the temperature $T(\theta_{i},\omega_{i})$ to the second order for minute quantities instead of truncating $\mathcal{O}\left(\Delta T\omega_{i},\omega_{i}^{2}\right)$ as in Eq. (61):
@@
The first term in Eq. (81) coincides with the averaged brake power $\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{M}}\Bigr{\rangle}$ of the engine ${\mathrm{M}}_{\mathrm{EG}}$ in system M, which can be confirmed by multiplying both sides of Eq. (32) by $N\omega_{\mathrm{m}}$ and averaging it over time. On the other hand, the second term in Eq. (81) agrees with the averaged brake power $\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{D}}\Bigr{\rangle}$ for system D. Therefore, Eq. (81) agrees with the averaged brake power corresponding to the conceptual model constructed in III. B. It can be seen that the power dissipated by the viscous friction of the conceptual model consists of the power dissipated by the viscous friction due to the rotational motion of the Stirling engines, $\sum_{i=1}^{N}\Gamma\Bigl{\langle}\omega_{i}^{2}\Bigr{\rangle}$, and the loss of the energy flow that the Stirling engines receive from the heat reservoirs, $\frac{\sigma^{2}}{G}\biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\biggr{\rangle}_{\theta}\left(\sum_{i=1}^{N}\Bigl{\langle}\omega_{i}^{2}\Bigr{\rangle}\right)$. The latter is the heat leakage due to nonquasi-static processes associated with the rotational motion of the Stirling engines, which depends on the magnitude of the coupling strength: the portion of the heat leakage that depends on the motion of the mean angle of the coupled Stirling engines, $\frac{\sigma^{2}}{G}\biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\biggr{\rangle}_{\theta}\left(\sum_{i=1}^{N}\Bigl{\langle}\omega_{\mathrm{m}}^{2}\Bigr{\rangle}\right)$, is kept at a constant value, while the remaining that depends on the relative motion, $\frac{\sigma^{2}}{G}\biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\biggr{\rangle}_{\theta}\left(\sum_{i=1}^{N}\biggl{\langle}\left(\omega_{\mathrm{d}}^{(i)}\right)^{2}\biggr{\rangle}\right)$, becomes smaller when the coupling strength is increased. Thus, the conceptual model constructed in III. B interprets the heat leakage due to nonquasi-static processes as part of energy dissipation due to viscous friction in the conceptual model, thereby allowing the original coupled system to be seperated into two isolated subsystems: a coupling strength-independent system M and a weakly-coupled system D.